# Loop quantum effects on direct detection prediction in two-scalar dark matter scenario Karim Ghorbani

### v1

@@ 4 Leading Quantum Corrections (LQC)
where the loop functions ${\cal H}_{1}$ and ${\cal H}_{2}$ are provided in Appendix C. The part of the amplitude which is proportional to $\kappa^{2}_{12}$ is dominant, besides that the box contribution itself is suppressed because of the two insertions of the quark-Higgs vertex.
@@ 8 Appendix B
where ${\cal C}_{q}=m_{q}\kappa_{22}/m^{2}_{h}$. The elastic scattering cross section of DM-nucleon can be obtained by replacing the quark current with nucleon current at cost of a proportionality factor. The final result we arrive at is a spin-independent (SI) DM-nucleon cross section
$\sigma^{\text{tree}}=\frac{F^{2}_{N}\mu^{2}_{N}}{\pi m^{2}_{2}}\,,$ (27)
in which the parameter $\mu_{N}$ is the reduced mass of the DM particle and the nucleon, and the parameter $F_{N}$ is connected to the scalar couplings, $f^{N}$ in the following way,
$F_{N}=\sum_{q=u,d,s}\frac{m_{N}}{m_{q}}{\cal C}_{q}f^{N}_{q}+\frac{2}{27}\sum_{q=c,b,t}\frac{m_{N}}{m_{q}}{\cal C}_{q}f^{N}_{g}=\Big{(}\frac{m_{N}}{m^{2}_{h}}\kappa_{22}\Big{)}f^{N}\,.$ (28)
Here $m_{N}$ is the nucleon mass, in our numerical calculations we set it equal to the proton mass, and also we take for proton, $f^{p}\sim 0.284$ [46].

### v2

@@ 4 Leading Quantum Corrections (LQC)
where the loop functions ${\cal H}_{1}$ and ${\cal H}_{2}$ are provided in Appendix C. The part of the amplitude which is proportional to $\kappa^{2}_{12}$ is dominant. Concerning light quarks in the nucleon, the box contribution is suppressed because of the two insertions of the quark-Higgs vertex.
5 Numerical Results
@@ 8 Appendix B
where ${\cal C}^{\text{tree}}_{q}=m_{q}\kappa_{22}/m^{2}_{h}$. The elastic scattering cross section of DM-nucleon can be obtained by replacing the quark current with nucleon current at cost of a proportionality factor. The final result we arrive at is a spin-independent (SI) DM-nucleon cross section
$\sigma^{\text{N}}=\frac{F^{2}_{N}\mu^{2}_{N}}{\pi m^{2}_{2}}\,,$ (30)
in which the parameter $\mu_{N}$ is the reduced mass of the DM particle and the nucleon, and the parameter $F_{N}$ is connected to the scalar couplings, $f^{N}$, in the following way,
$F_{N}=\sum_{q=u,d,s}\frac{m_{N}}{m_{q}}{\cal C}_{q}f^{N}_{q}+\frac{2}{27}\sum_{q=c,b,t}\frac{m_{N}}{m_{q}}{\cal C}_{q}f^{N}_{g}\,,$ (31)
where at tree level, we have
$F_{N}=\Big{(}\frac{m_{N}}{m^{2}_{h}}\kappa_{22}\Big{)}f^{N}\,.$ (32)
Here $m_{N}$ is the nucleon mass, in our numerical calculations we set it equal to the proton mass, and also we take for proton, $f^{p}\sim 0.284$ [49], given that $f^{N}_{g}=1-\sum_{q=u,d,s}f^{N}_{q}$. The second term in Eq. 31 arises from effective DM-gluon interactions which is obtained based on a relation between $G^{a}_{\mu\nu}G^{a\mu\nu}$ and heavy quark current, $\bar{Q}Q$, [55, 33]
$m_{Q}\bar{Q}Q=-\frac{\alpha_{s}}{12\pi}G^{a}_{\mu\nu}G^{a\mu\nu}\,.$ (33)
When we go beyond tree level then contributions from triangle and box diagrams (including effective DM-gluon interaction) add extra terms to ${\cal C}_{q}$ as ${\cal C}_{q}={\cal C}^{\text{tree}}_{q}+{\cal C}^{\text{triangle}}_{q}+{\cal C}^{\text{Box}}_{q}$. As pointed out in [33] a complete treatment of DM-gluon interaction via box diagrams requires full two-loop computations which is beyond the scope of the present work.